Two buses leave the station at the same time. One travels east at 55 miles per hour. The second travels west at 60 miles per hour. In how many hours will the two buses be 460 miles apart?
Answer by ewatrrr(24785)


Hi,
Let x represent time each traveled. d = r*t
55mph*t + 60mph*t = 460mph
t(55mph + 60mph) = 460mph
t = 460/115
t = 4hr
checking our answer
4*55 + 4*60 = 220 + 240 = 460